# Secure Human Oversight of AI: Exploring the Attack Surface of Human Oversight

### v1

@@ 4.7 Training of Human Oversight Personnel
Training can also address resilience against coercion, preparing personnel to respond to threats or undue pressure. This includes protocols for escalating coercion attempts, access to secure whistleblowing channels, and preparation to maintain effective human oversight behavior under stress. Anti-bribery training can teach personnel to recognize undue influence. Clear organizational policies defining conflicts of interest and mandatory disclosure mechanisms further reduce the risk of misaligned intentions. Training of human oversight personal counters social engineering, coercion, and bribery.
Hardening strategy: Organizations should implement recurring, context-specific training programs for human oversight personnel. By institutionalizing such training and embedding it in the organization’s security culture, human oversight personnel can develop the awareness, self-efficacy, and procedural support needed to counteract manipulation attempts, thereby contributing to secure human oversight.
@@ 5 Discussion
Research on human oversight of AI has primarily focused on increasing its effectiveness [1, 2, 6, 4]. In this paper, we highlight the importance of complementing this focus with a security-related perspective on human oversight. This perspective is critical because various attack vectors can undermine human oversight effectiveness, a risk that will grow as human oversight becomes more widely implemented as part of the safety, security, and accountability architecture of AI operations. Drawing on cybersecurity research [33, 44], we have introduced various attack vectors targeting human oversight, along with corresponding hardening strategies, to support the development of secure human oversight of AI.
Understanding and designing the technical factors, selecting and training human oversight personnel, and shaping working conditions to foster effective human oversight remains an ongoing challenge for research, practice, and AI governance [1, 2, 95]. Our analyses indicate that this challenge will intensify as human oversight becomes standard practice in high-risk AI use cases. Malicious actors targeting the attack surface of human oversight can compromise the AI system, the communication between humans and AI, or the oversight personnel themselves, aiming to undermine epistemic access, causal power, self-control, and fitting intentions as the core requirements for effective oversight [1]. Consequently, such attacks may increase risks in AI operations by reducing the effectiveness of oversight, disabling oversight entirely, or manipulating oversight personnel’s intentions to exacerbate risks. The attack vector list we present is not exhaustive; future research should explore novel attack vectors specific to human oversight. A potential future scenario involves the deployment of malicious AI agents designed to deceive oversight [70, 69]. Once an AI system models the oversight mechanism, it may attempt to circumvent the restrictions imposed by human oversight, further expanding the attack surface and associated risks.
@@
Given these risks, it is essential to consider how to enhance the security of human oversight. To this end, we presented various hardening strategies drawn from the broader cybersecurity literature [52, 53, 54]. These strategies must be actively applied and adapted to the context of human oversight of AI. Our paper contextualizes these strategies for human oversight, representing a crucial first step toward securing oversight architectures against malicious attacks. Future research needs to evaluate the effectiveness of these hardening strategies and could develop additional strategies tailored specifically to human oversight. Beyond technical hardening, future work could also explore the application of amplified oversight. The core principle of amplified oversight is to employ AI itself to systematically reduce complexity to a level that enables human oversight personnel to fulfill their duties effectively [96]. In future, such approaches to support human oversight may even exceed human capabilities and may become a further viable target for malicious actors to attack.
While this paper focuses on malicious actors targeting human oversight, not all risks emerge from malicious actors. Additional risks can for example arise from counterproductive behavior within organizations [97]. Counterproductive behavior refers to intentional or unintentional actions by employees that go against organizational goal, such as neglecting oversight duties, resisting established procedures, or even deliberately undermining AI operations, which can erode the effectiveness of human oversight. Such behavior may stem from factors like dissatisfaction with working conditions, misaligned incentives, organizational conflict, or lack of recognition [97]. Such counterproductive behavior inside organizations may need closer examination in the future to ensure secure human oversight. Such behavior can undermine the core requirements of effective oversight and may promote active damage to the organization or passive under-performance, thereby introducing risks to human oversight of AI. Although these behaviors are not attacks by malicious actors, they may make attacks by external actors more effective, for example, if an attacker targets a dissatisfied employee. Future works could shed light on those risk factors and suitable hardening strategies such as organizational structures, incentive systems, and thoughtful work design [29]. Well-designed security concepts, policies and work conditions can all together mitigate security risks and prevent counterproductive behavior by ensuring that oversight personnel’s goals remain aligned with organizational objectives.
@@ 6 Conclusion
Research on human oversight has primarily focused on its effectiveness – and rightfully so, given the complexity of designing oversight processes that genuinely mitigate risks in AI operations [1, 3, 2]. Adding to this complexity, it will become increasingly important to design for secure human oversight as oversight becomes an integral component of the safety, security, and accountability infrastructure of AI operations. We hope this paper serves as a starting point for research that analyzes the attack surface of human oversight and develops effective hardening strategies to prevent malicious actors from exploiting oversight as a gateway to compromise AI safety.

### v2

@@ 3.3.5 Training of Human Oversight Personnel
It can also address resilience against coercion, preparing personnel to respond to threats or undue pressure. This includes protocols for escalating coercion attempts, access to secure whistleblowing channels, and preparation to maintain effective human oversight under stress. Anti-bribery training can teach recognizing undue influence. Clear organizational policies defining conflicts of interest and mandatory disclosure mechanisms further reduce the risk of misaligned intentions. Training of human oversight personal counters social engineering, coercion, and bribery.
3.3.6 Red Teaming
The term red team refers to an independent group that investigates an organization’s IT systems to identify vulnerabilities and security flaws [49]. The red team simulates the behavior of malicious actors, employing all methods that real attackers might use. Its goal is to uncover as many exploitable security issues within the tested system as possible. Continuous red teaming can help to stay on top of emerging security flaws after the initial threat modeling of an application. Software systems are dynamic even in deployment and, thus, require constant monitoring.
@@ 4 Discussion
Research on human oversight of AI has primarily focused on increasing its effectiveness [72, 25, 18, 67]. In this paper, we highlight the importance of complementing this focus with a security-related perspective on human oversight. This perspective is critical because digital systems are increasingly pressured by cyber criminals and weaknesses in single components can compromise whole IT infrastructures, a risk that will also apply to human oversight as it becomes more widely implemented as part of the safety, security, and accountability architecture of AI operations. Drawing on cybersecurity research [65, 15], we have introduced threat modeling for human oversight. We designed a DFD for the human oversight application, identified assets that need to be protected, and explored vulnerabilities of human oversight using STRIDE. Furthermore, we introduced a minimal set of mitigation strategies to secure human oversight.
Understanding and designing the technical factors, selecting and training human oversight personnel, and shaping working conditions to foster effective human oversight remains an ongoing challenge for research, practice, and AI governance [72, 25, 44]. Our analyses indicate that this challenge will intensify as human oversight becomes standard practice in high-risk AI use cases. By analyzing an abstracted application-level representation of human oversight using structured threat modeling, we unveiled vulnerabilities that can be exploited by malicious actors to attack key assets, including the requirements of effective human oversight [72].
Consequently, such attacks may increase risks in AI operations by reducing the effectiveness of oversight, disabling oversight entirely, or manipulating oversight personnel’s intentions to exacerbate risks. We were able to identify a list of security risk using the STRIDE model. However, this list can only serve as a starting point for real applications of human oversight. In practice, our baseline has to be extended with the concrete software and hardware components that are integrated into the human oversight application. These components can open an application to new attack vectors. Furthermore, we included threats posed by misaligned AI agents into our investigation. With growing knowledge about AI agents in practice, new ways possibly emerge how AI agents can deceive oversight [51, 7].
While we were using STRIDE to analyze our theoretical human oversight application, additional options are available for the secure implementation of human oversight. Once human oversight is operational, the analysis with STRIDE should be supplemented with operational threat modeling using frameworks like cyber kill chain [79], MITRE ATT&CK [4], or unified kill chain [59]. This allows for an even more comprehensive and specific analysis of the operational aspects of a human oversight application. Another endeavor that should be considered by practitioners when performing real-world threat modeling is the quantitative analysis of threats. This means determining for each threat the ease of exploitation and the damage potential of an exploitation. Such analysis can help to prioritize hardening strategies.
Identifying risk is only one side of the cybersecurity coin. Appropriate hardening strategies have to be put into place to secure applications. We introduce several hardening strategies that can be used to mitigate the identified vulnerabilities. These strategies are based on the cybersecurity literature [62, 61, 9]. However, our provided list has to be understood as a minimal set of measures. Future research needs to identify additional measures and evaluate the effectiveness of common hardening strategies for human oversight.
@@
Beyond technical hardening, future work could also explore the application of amplified oversight. The core principle of amplified oversight is to employ AI itself to systematically reduce complexity to a level that enables human oversight personnel to fulfill their duties effectively [66]. In future, such approaches to support human oversight may even exceed human capabilities and may become a further viable target for malicious actors to attack.
@@ 5 Conclusion
Research on human oversight has primarily focused on its effectiveness – and rightfully so, given the complexity of designing oversight processes that genuinely mitigate risks in AI operations [72, 45, 25]. Adding to this complexity, it will become increasingly important to design for secure human oversight as oversight becomes an integral component of the safety, security, and accountability infrastructure of AI operations. We hope this paper serves as a starting point for research that analyzes the cybersecurity risk of human oversight. Additionally, we hope our paper serves as a guideline for researchers and technicians that are task to implement human oversight in a real-world system.